OpenStudy (anonymous):

0.5000 kg of water at 35.00 degrees Celsius is cooled, with the removal of 6.300 E4 J of heat. What is the final temperature of the water? Specific heat capacity of water is 4186 J/(kg Co). Remember to identity all of your data, write the equation, and show your work.

OpenStudy (anonymous):

and also How would you explain the concept of an expanding universe to a fellow scientist?

OpenStudy (anonymous):

I would greatly appreciate some help!!! Thank you!!!!! Medals will be awarded!

OpenStudy (matt101):

You can solve this question using the following equation: \[q=mc \Delta T\] Where q is the total heat absorbed/released, m is the mass, c is the specific heat, and ΔT is the change in temperature.

OpenStudy (shamim):

The above equation is right !!!!!

OpenStudy (shamim):

Given q=6.3*E4 joules c=4186j/kg/c Theta1=35degree celcius Theta2=?

OpenStudy (shamim):

We can rewrite the above equation as q=mc*(theta1-theta2) Theta2=?